Home‎ > ‎Crew 132 - Mars Society‎ > ‎

1224-CommandersReport

Commander Report
Nick Orenstein

Commander's Log, Mars Date 1321224.4

Leave only footprints, take only pictures. Today, my footprints touched Mars. I led the away team on our maiden journey to the surface of the Red Planet. Our habitat must be located close to the North Polar region. Much of the ground was covered in crystalized water ice. Indeed, our dreams of a white Christmas EVA came true.
This his eve of a classical Earth religious holiday has inspired Crew 132 to discuss their diverse range of beliefs. Heritages range from the spectrum of Earth religions. Some crew members agree that Earthlings, and especially those of spacefaring nature, are embarking on a new historical age. Organized religion is transitioning to individual interpretations of spirit. Ancient cultures' cosmovision is being exposed through the modern study of archeoastronomy. I cherish my family's religious traditions and joyfully celebrate with others for theirs.
When starting this log, at the crew's urging, I donned a fuzzy red and white hat and became St. Nick. Before finishing, I was interrupted to climb a crawl space to fix the hot water heater. Little does the crew know that Santa is coming back down the chimney with a special gift from Earth for these nice boys and girls. 
Ho ho ho, Merry Christmas.

___________________________


Commander Check-In Report
Nick Orenstein

Crew Physical Status: 
No reported physical ailments so there will be no HSO report. EVA team exerted themselves but recovered quickly.

Eva Depart Time: 10:15
Eva Return Time: 12:50
EVA Highlights : An away team of four journeyed around the Hab and to a nearby hill. This was our crew's first EVA so we used it to debug our EVA procedure and equipment. Helmets were very foggy. Radios were faulty. The team enjoyed the chance to explore the surface.

Engineering/Hab Maintenance:
Yesterday we did not actually turn on the water heater. We turned on an A/C unit. Today we found the water heater. It did not turn on. We debugged the problem which was that the Sci 1 breaker had switched off. Now the hot water heater is on.

Report Transmission Schedule:
Commander Check-In
Commander Report
Engineering Report
EVA Report
Future EVA Plan
Daily Photos
Journalist Report

Plans for Tomorrow:
First biology/geology EVA. Cut and assemble GreenHab hydroponics PVC and hose. 

Inventory: 
- Hung new 2 solar-powered "Little Sun" lights in the GreenHab donated by artist Olafur Eliasson. 
- Two of the charging cables (#5 and #2) for new EVA backpack fans is operable.
- All of the EVA radio headsets are inoperable.
- The clock on the microwave in the Hab strangely loses time. Perhaps it is set on a day/sol delay to Earth time and not local Mars time.

Support Requests:
Updated Wifi settings are included in the Engineering Report. Tell Judd thanks for his email about HAL. Access to HAL is still not 100% but now we think we understand how to fix things. Will attempt tomorrow and report back. Still submitting Comms via email.

Miscellaneous:
Christmas Eve on Mars.